• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Yi Zhang

    Log entry time 08:05:37 on November 4,2008

    Entry number 246601

    keyword=Owl Shift summary

    SL: Yi Zhang
    TO: Wooyoung Kim
    3rd: Huaibo Ding

    Summary:
    longitudinal elastic
    2771 2M
    2772 2M
    2773 2M
    2774 1.8M
    Transverse+ delta run
    2777 haven't recorded, supposed to be 0.6M
    2779 2M
    2780 1.48M
    2781 1.1M
    2783 1.5M
    2786 0.6M
    2787 1M

    Activities:
    00:00 Shift start. Everything goes smoothly.
    00:17 Stop run#2771, It has 2M events
    00:18 Start run#2772.
    00:58 Stop run#2772. It has 2M events.
    01:05 Finish a NMR measure in pick-up coil position. We got ~36% polarization, So continue longitudinal elastic run.
    01:10 Start run#2773.
    01:44 Stop run#2773. It has
    01:44 Start run#2774
    02:22 Stop run#2774 because beam trip for 5min. It has 1.8M events.
    02:23 Do another NMR, get a polarization about 30%. Jin suggests go transvers pumping run.
    02:32 During rotating the field, Xin comes and ask for a checking run for BB chamber. the run number is 2776(2775 is junk.) Now Jin is working on field rotating (to transverse+).
    02:37 2776 stoped. Xin is analysing it now. Later he finds it's still in bad shape.
    02:37 Jin asks TO move target from pick-up coil position (for NMR) to optics position (for NMR calibration). Then he asks to turn on the transverse laser(1,3,4,5,6).
    02:44 Jin ask TO lower the target to He3 cell position, Then he do a NMR again. From 2 NMR, we know the ratio between 2 NMR signal in 2 different position (PU coil & optics).
    02:47 We change kinmatic of LHRS to 850MeV/C and ask for beam back.
    02:55 LHRS is ready. We are waiting for stable beam now
    03:04 We start new delta run(transverse pumping). The run number is
    2777. But later we find the dead time is -5, there should be something wrong.
    03:15 Stop run#2777. Because beam trip for 5+ min.
    03:16 Jin does another NMR for checking we are pumping the cell, not depolarizing. The result shows the signal is rising.
    03:20 MCC calls us for a long beam trip(about 0.5 to 1 hour as he says).
    03:22 Call MCC to move target to optics without hole, just for safe.
    04:13 MCC call us to deliver beam. We put BeO target first to have a check.
    04:17 Start Run#2779
    04:57 Stop Run#2779. It has 2M events.
    04:58 Start Run#2780
    05:24 Stop Run#2780 for start Happex daq. It has 1.48M events.
    05:26 Start Run#2781(LHRS&BB) and Run#30364(HAPPEX)
    05:33 Stop run#30364 as Bryan told me. Then restart coda for happex daq and start a new run#30366 @ 05:38
    05:50 Stop Run#2781&30366, call MCC to stop beam, have a NMR measurement in He3 target position.
    05:56 Start Run#2783&30367 (2782 is junk)
    06:27 Stop Run#2783#30367 for BB HV alarm. No flashing, and the rates looks nomally.
    06:28 coda reboot as find some problem on the halog of run#2783 starting
    06:39 Start Run#2786 (2785 is junk, 2784 is for coda checking), after talking with Kalyan about BB HV status. Now I'm sure the HV of BB works fine.
    06:51 Stop Run#2786, finding there is a problem in g0 helicity start from run#2779. Page Bryan again.
    06:55 Start Run#2787&30370, without getting reply from Bryan.
    07:21 Stop Run#2787&30370. Run 2787 has 1M events. Then Jin starts manually flit target spin.
    07:37 Jin has done. Ask beam back. Start taking data. The run number is 2788&30371. By the way, after spin flip, we do another NMR, which tells us the polarization is 50%